OpenStudy (anonymous):

express [1/5(x+2)]+[13/5(2x-1)] in form of a(1/x)+b(1/x)^2+c(1/x)^3+... (up to 3 term) where a,b,c is constant

OpenStudy (anonymous):

\[1\div[5(x+2)]+13 \div[5(2x-1)]\]

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

myininaya never sleeps

myininaya (myininaya):

satellite do you know how to do this? i don't think it is possible is it?

myininaya (myininaya):

wait you have +... so you are talking about series expansion

OpenStudy (anonymous):

addition of power series

OpenStudy (anonymous):

ya... this special question.... actually my teacher give me question that express in (1+x)^n =1+nx+[(n)(n-1)/2! ](x)^2+[n(n-1)(n-2)/3!](x)^3+...

OpenStudy (anonymous):

factor out the 1/5 first probably

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then get series for \[\frac{1}{x+2}\] i would cheat but the derivatives are easy

OpenStudy (anonymous):

(1+x)^n =1+nx+[(n)(n-1)/2! ](x)^2+[n(n-1)(n-2)/3!](x)^3+... actually i use this equation to do

OpenStudy (anonymous):

\[f'=-(x+2)^{-2}\] \[f''=2(x+2)^{-3}\] \[f^{(3)}=-6(x+2)^{-4}\] etc

OpenStudy (anonymous):

ok so coefficients are easy enough. alternating series

OpenStudy (anonymous):

looks like the factorials are going to cancel and you will get something in terms of \[2^n\]right?

OpenStudy (anonymous):

\[\frac{1}{2}-\frac{x}{4}+\frac{x^2}{8}-...\]

OpenStudy (anonymous):

of course that is just for \[\frac{1}{x+2}\] then you have to do it again for \[\frac{13}{2x-1}\] then add

OpenStudy (anonymous):

you think it would be easier to add first expand second?

myininaya (myininaya):

i think we should do the expansions and then add

myininaya (myininaya):

but either way

OpenStudy (anonymous):

yeah you are right i just added via wolfram http://www.wolframalpha.com/input/?i=1%2F%285%28x%2B2%29%29%2B13%2F%285%282x-1%29%29

OpenStudy (anonymous):

my teacher teach to expand the first series and the second, then sum up the series

OpenStudy (anonymous):

ooooh look it even provides the expansion. but i guess the question is how to get there. gotta run

myininaya (myininaya):

later and good job

OpenStudy (anonymous):

ok... maybe i can try to figure out

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

ok... i get it... first factories 1/x first then i can get... in term a(1/x)+b(1/x)^2+c(1/x)^+...